OpenStudy (anonymous):

Which system of linear equations has no solution? A. 2x − y = −1 2x + y = 2 B. 4x − y = 3 3x − 2y = 3 C. −2x + 4y = −16 4x − 2y = −1 D. 2x − y = −1 −2x + y = −1

OpenStudy (anonymous):

my choice would be A. but I'm not sure.

OpenStudy (eric_d):

2x-y=1 2x-1=y 2x+y=2 y=-2x+2 2x-1=-2x+2 4x=3 x=3/4 y=2(3/4)-1 y= It can't be A

OpenStudy (eric_d):

Do solving other choices.. if u can't get value of x,y ... that'll be ur answer

OpenStudy (eric_d):

Try solving*